What is the equation to solve this word problem: A bank deposit grew to $2000 in 10 months paying 8% interest. What is the principal? I tried:
2000=x * .08 * (10/12) but not sure if this is correct and don't understand next steps.

Answer by robertb(5830)

If it is simple interest, then your equation is B = P(1+rt).
2000 = x(1+0.08*(10/12)),
2000 = x(1.066666666...),
x = 1875.